Congratulations! You are hired in a business organization. In this company the market research department that you are the member recommended to manufacture and market a promising new product (x). After extensive surveys, the research department supported the recommendation with the demand function:



Dx = f (p) = 60 – 3 P



Where Dx is quantity demanded at price P during a month. Financial department has brought the cost function and is given by



C(x) = 72 + 6x



Now the team leader of the market research department ordered you to manipulate the following problems.



Express revenue as a function of X



Expresses profit as a function of X

1
Expert's answer
2022-02-06T14:23:11-0500

Solution:

Express revenue a a function of X:

Revenue function = P "\times" Q

TR = P "\times" Q

x = 60 – 3P

P = 20 – "\frac{x}{3}"

TR = (20 – "\frac{x}{3}") "\times" X = 20X – "\frac{X^{2}}{3}"

Total Revenue function = 20X – "\frac{X^{2}}{3}"

 

Expree profit as a function of x:

Profit = TR – TC

Profit = (20X – X2/3) – (72 + 6x)